8. Determinants of intrinsic value Aa Aa Which of the following is a determinant of a firms cost of capital? Cost of debt O Education of managers O Sales revenues The following table presents some of the current years financial data for the Happy Giraffe Construction Company. Complete the table by computing Happy Giraffes free cash flows and fundamental value (rounded to the nearest whole dollar) based solely on this years information: Value $400,000 $45,000 $20,000 15% $335,000 $2,233,333 Variable Sales revenue Operating costs and taxes Required investments in operating capital Weighted average cost of capital Annual free cash flow Fundamental value Assuming all other things being equal, if Happy Giraffes costs of equity decreases, the firms will

Please answer
Answer #1

Cost of debt is a determinant of capital.

Sales Revenue = $400,000

Less: Operating Costs = $45,000

Less; Cost of Capital = 20,000*15% = $3,000

Annual Free cash flow = $352,000

Funadamental Value = $352,000/0.15 = $2,346,667